OpenStudy (anonymous):

solve system y=x+2 and y=-4x+12

OpenStudy (anonymous):

5x=10... get the ans

OpenStudy (anonymous):

Hi The first step is to make them equal to each other. This is because y =y. Therefore set out as x + 2 = -4x +12 gather like terms Get all the x to one side, and then numbers to the other x + 2 = -4x +12 x+4x +2 = +4x +12 5x +2 =12 5x -2 = 12 - 2 5x = 10 solve for x by dividing by 5, x = 2 Now sub in this x value we have found into the original equation x + 2 = -4x +12 2 + 2 = -4(2) +12 4 = -8 +12 4 = 4 y=y y = 4 Hope it helps let me know good luck